Mathematics 133 – Intensive Calculus for Science 1 Quiz 9 – November 18, 2005

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. Calculators allowed on this quiz.

A) Find the maximum and minimum values of f(x) = x^{2}−4x+2 on the interval 0 ≤ x ≤ 5.

B) Find the maximum area for a rectangular garden enclosed by 2000 feet of fencing, if one side of the rectangle is formed by part of an existing long stone wall.